# version 0.6

### v3

@@ 11.6 Exercises
11.2. Prove that the class $\prod^{LIN}$ $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters.
Chapter 12 Kalman Filtering
The previous chapter introduces the concept of open-loop optimal linear predictors and how to efficiently learn them using convex relaxation and online gradient methods. However, a good predictor should exploit not only its initial knowledge, but also information as it is being observed and learn from it. In an online prediction setting, this amounts to using the online observations to further refine future predictions.
This is the topic of this chapter: how to use feedback from the ongoing prediction task to improve future predictions. Optimal closed-loop predictors, that use feedback from observations, are significantly more powerful and harder to analyze. We explore Kalman filters, prove their optimality, as well as convex relaxations that fit into the paradigm of online convex optimization.
@@ 12.1 Observable Systems
In the discussion below, when considering the infinite horizon setting, prediction makes sense if our observation remains bounded. The condition for bounded observations is related to controllability as defined in Section 4.3.
We first state the definition of observability for general dynamical systems below. In absence of noise and control inputs, every dynamical system can be viewed as mapping from an initial state to an infinite sequence of observations. Observability asserts that this mapping is injective.
@@
A partially-observed dynamical system is said to be observable if and only if every sequence of observations produced by the deterministic-equivalent system uniquely determines the initial hidden state.
For linear systems, checking observability can be boiled to a rank condition on the Kalman observability matrix, which is defined as
$K^{\prime}_{k}(A,C)=\begin{bmatrix}C\\ CA^{2}\\ \dots\\ CA^{k-1}\end{bmatrix}.$
We ask the reader to provide a proof of the following result in Exercise 12.5.
@@
A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is said to be observable if $\mathop{\mbox{\rm rank}}(K^{\prime}_{d_{x}}(A,C))=d_{x}$.
12.2 Closed Loop Optimal Predictors
The previous class of predictors is open-loop, namely it only takes into account the inputs to the system, which can potentially be computed ahead of time, and not the response of the system to the controls and perturbations. A closed-loop prediction rule takes into account the observations, and can potentially react to the perturbations and observation noise.
The most famous example of such a prediction rule is the Kalman filter, which we now define. The defining characteristic of the Kalman filter is that it produces the best linear predictor in the least-squares sense when the perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are zero-mean and i.i.d., and when the systems matrices $(A,B,C)$ and the perturbations’ covariances $(\Sigma_{w},\Sigma_{v})$ are known and defined as
$\Sigma_{x}\triangleq\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
So far we have focussed on predicting observations, i.e., ${\mathbf{y}}_{t}$’s. On the other hand, the next theorem produces optimal estimators $\hat{\mathbf{x}}_{t}$ of the hidden state $\mathbf{x}_{t}$ given past observations $\mathbf{y}_{1:t-1}$ and controls $\mathbf{u}_{1:t-1}$. This is motivated by the fact that in the stochastic case, the optimal least-squares predictor can be written as $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$. In Exercise 12.5, we ask the reader to supply a proof of this statement.
@@ 12.2 Closed Loop Optimal Predictors
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
Proof.
$\Sigma_{t}\triangleq\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
We begin with an observation about least-squares predictors, which we ask the reader to prove in Exercise 12.5.
@@
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{LIN}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 12.5).
@@
$A^{*}\triangleq\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
@@
$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{LIN}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
@@
$\hat{\mathbf{x}}_{t+1}=A\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}\left(\Sigma_{y}+C\Sigma_{t}C^{\top}\right)^{-1}}_{\triangleq L_{t}}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),$
@@
$\Sigma_{t+1}$ $\triangleq\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
@@
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. Exercise 12.5 asks for a proof of the next lemma.
@@
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{LIN}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@ 12.3 Bayes-optimality under Gaussian Noise
for some $\Sigma_{t}\in{\mathbb{R}}^{d_{x}\times d_{x}}$, for which we provide a recursive formula towards the end of the proof. We make note of the following lemma for Gaussian distributions, whose proof we ask the reader to complete in Exercise 12.5.
@@
$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\triangleq L_{t}}\mathbf{y}_{t}.$
@@ 12.5 Exercises
12.1. Prove Lemma 12.5.
12.2. Using Lemma 12.5, prove Lemma 12.4.
12.3. Prove Lemma 12.6.
12.4. Prove Lemma 12.8.
12.5. Prove Theorem 12.2.
@@
Chapter 13 Spectral Filtering
In the previous chapter we have showed how to learn an unknown linear dynamical system with perturbations and noise using convex relaxation and online learning. Instead of learning the system matrices directly, we have learned an alternate class of predictors, namely the class of linear predictors over past observations and controls, that approximates the optimal predictor in certain cases.
Despite the clean and simple formulation of this result, there remains an unsatisfying issue with the previous chapter. The number of parameters required to learn the optimal open loop predictor in the class $\prod_{\star}$ depends on the measure of stability of the underlying system, which is potentially large and unknown.
In this section we describe an elegant alternative method for learning a linear dynamical system that removes this dependence on the stability constant of the system. For the sake of exposition, we restrict ourselves only to symmetric system matrices $A$. This is a significant restriction, which can be removed, albeit using a significantly more sophisticated approach which is referred to in the bibliographic section.

### v4

@@ 9.6 Bibliographic Remarks
For a comprehensive treatment of online learning in games and online convex optimization, see the texts [CBL06, Haz16].
In the statistical noise setting, a natural approach for learning is to perform system identification and then use the identified system for prediction. Such an approach was taken by [SMT+18, SF20, SR19]. The work of [GLS+20] extends identification-based techniques to adversarial noise and marginally stable systems.
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18]. In that paper, it was shown how to learn partially observable, and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
@@ 9.7 Exercises
9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero-mean noise.
9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
9.4. Give precise regret bounds for learning the class $\prod^{L}_{h,k,\kappa}$ and $\prod^{\star}_{h}$ as a function of the natural class parameters $h,k,\kappa$ and any other relevant parameter.
9.5. Consider the square loss in algorithm 6. Give a Newton-method based algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
9.6. Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time varying linear dynamical systems. Research the notion of adaptive regret from [Haz16], and reason about conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time varying linear dynamical systems.
Chapter 10 Kalman Filtering
State estimation plays a central role in control and prediction tasks, where an agent seeks to infer latent states from noisy observations. The Kalman filter is a classical and widely used method for solving this problem, providing optimal state estimates under the assumption of linear system dynamics and Gaussian noise.
A key feature of the Kalman filter—and a fundamental limitation—is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model. These matrices encode how the system evolves over time and how observations relate to the hidden state. While the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free, open-loop alternative by directly learning optimal linear predictors from data. Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems. However, these approaches do not incorporate feedback, meaning they operate in an open-loop setting, predicting future states without correcting for new observations.
This chapter moves beyond open-loop prediction to closed-loop state estimation, where past observations continuously refine future predictions. The Kalman filter accomplishes this by recursively updating estimates based on incoming measurements, achieving Bayes-optimal performance under Gaussian noise. This structured feedback mechanism allows for more accurate and robust estimation compared to open-loop methods, particularly in dynamic and uncertain environments.
Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter. This approach retains the feedback mechanism of Kalman filtering while removing the need for explicit system identification, making it applicable in adversarial or uncertain environments. By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
We provide a rigorous introduction to Kalman filtering, including its derivation, optimality properties, and computational aspects. We also discuss its limitations and contrast it with learning-based alternatives, setting the stage for spectral filtering techniques in the next chapter, which further generalize state estimation in unknown dynamical systems.
@@ 10.1 Observable Systems
Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its outputs over time. In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
A general dynamical system maps an initial state to an infinite sequence of observations. If this mapping is injective—meaning that each initial state corresponds to a unique sequence of observations—then the system is said to be observable. Formally:
@@
A partially observed dynamical system is observable if and only if every sequence of observations uniquely determines the initial hidden state.
A simple example of an unobservable system is given as the following variant of the double integrator from section 2.1.2. Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given as follows
[…]
This system allows only the velocity of the object to be observed, while the position remains hidden. It can be seen that this setting is unobservable, as the initial position cannot be determined based on the velocities and forces applied to the system alone. Indeed, this is a special case of the Galilean Relativity principle of Newtonian mechanics, which states that the laws of physics are the same in all inertial reference frames.
For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by:
$\hat{K}(A,C)=\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}.$
[…]
@@
A partially-observed dynamical system specified by system matrcies $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
10.2 Optimality of the Kalman Filter
10.2.1 The Power of Linear Predictors
One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed loop policies. We prove this by proving two facts:
  1. 1.
    The class of linear policies $\varepsilon$-approximates the class of Kalman filters, in the sense of Definition 9.1.
  2. 2.
    The class of Kalman filters is optimal for closed loop policies.
The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed loop optimal predictor in hindsight. If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter. We prove this formally at the end of this section.
10.2.2 Closed Loop Optimal Predictors
Open-loop predictors, such as those introduced in the previous chapter, leverage system inputs and initial conditions to make predictions. However, they do not incorporate feedback from observations, making them susceptible to modeling errors and disturbances.
Closed-loop prediction rules address this limitation by continuously adjusting estimates based on new observations. The most well-known example of such a predictor is the Kalman filter, which produces the best linear predictor in a least-squares sense under the assumption that process and observation noise are Gaussian with known covariances.
Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices:
$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
So far, we have focused on predicting observations, i.e., $\mathbf{y}_{t}$. However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$. This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
$\mathbf{y}_{t}=C\mathbf{x}_{t}.$
In Exercise 10.6, we ask the reader to supply a proof of this statement.
Based upon this observation, the following theorem is the essence of the Kalman filter, which recovers estimates for states as well as linear projection matrices for optimal prediction of the next observation.
@@ 10.2.2 Closed Loop Optimal Predictors
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
To prove this theorem, we require some basic tools from linear regression which we state next.
10.2.3 Tools from Linear Regression
We begin with an observation about least-squares predictors, which we ask the reader to prove in Exercise 10.6.
@@ 10.2.3 Tools from Linear Regression
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
@@
$A^{*}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
@@
10.2.4 Analysis
We are now ready to prove the main property of the Kalman filter below.
Proof of Theorem 10.3.
$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
@@ 10.2.4 Analysis
$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{L}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
@@
$\hat{\mathbf{x}}_{t+1}=A\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}\left(\Sigma_{y}+C\Sigma_{t}C^{\top}\right)^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),$
@@
$\Sigma_{t+1}$ =$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
@@ 10.2.5 Infinite Horizon Kalman Filtering
10.2.6 Linear Predictors and the Kalman Filter
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is fomally stated in the following Lemma, whose proof is left as Exercise 10.6.
@@ 10.2.6 Linear Predictors and the Kalman Filter
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@ 10.3 Bayes-Optimality under Gaussian Noise
for some $\Sigma_{t}\in{\mathbb{R}}^{d_{\mathbf{x}}\times d_{\mathbf{x}}}$, for which we provide a recursive formula towards the end of the proof. We make note of the following lemma for Gaussian distributions, whose proof we ask the reader to complete in Exercise 10.6.
@@
$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
@@
10.4 Conclusion
In this chapter, we explored the Kalman filter as a fundamental tool for state estimation in dynamical systems, particularly in the presence of noise and uncertainty. We established its optimality in the least-squares sense and examined its connections to online convex optimization, reinforcing its relevance in both classical control and modern learning-based paradigms.
Key insights from this chapter include the role of observability in ensuring accurate state estimation, the derivation of the Kalman filter update equations, and the recursive structure that allows for efficient real-time implementation. We also presented the filter’s steady-state behavior, particularly in the infinite-horizon setting, where it converges to a solution governed by the Riccati equation.
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications, ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters), and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
@@ 10.5 Bibliographic Remarks
State estimation is a fundamental problem in control and signal processing, with numerous approaches beyond the classical Kalman filter. Alternative methods include:
  • Nonlinear Filtering Extensions.
    The Kalman filter assumes linear system dynamics and Gaussian noise, limiting its effectiveness in nonlinear settings. The Extended Kalman Filter (see e.g. [UA21]) addresses this by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, the linearization can introduce significant errors in highly nonlinear systems. Notable extensions using more sophisiticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
  • Particle Filtering (Sequential Monte Carlo Methods)
    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
  • Adversarial & Robust Filtering
    Robust filtering methods seek to provide guarantees under adversarial noise. The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-square error, making it suitable for applications with unknown disturbances.
While the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings. In the next chapter we will study an alternative using a technique which circumvents the need for state estimation altogether, and has stronger guarantees for sequential prediction.
@@ 10.6 Exercises
10.1. In this exercise we prove Theorem 10.2 in two parts:
  1. 1.
    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state. Hence the initial state is not uniquely determind.
  2. 2.
    Conversely, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
10.2. Prove Lemma 10.5, which states the following. Let $X,Y$ be possibly correlated random variables. Then
$A^{*}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\operatorname*{\arg\min}_{A}\mathbb{E}\|X-AY\|^{2}=\mathbb{E}\left[XY^{\top}\right]\left(\mathbb{E}\left[YY^{\top}\right]\right)^{-1}.$
Furthermore, if $\hat{X}=A^{*}Y$, then $\mathbb{E}(X-\hat{X})Y^{\top}=0$.
10.3. Using Lemma 10.5, prove Lemma 10.4, which states the following. Let $X,Y,Z$ be matrix-valued random variables. Let
If $\mathbb{E}\left[YZ^{\top}\right]=0$, then
$\mathbb{E}\|X-A^{*}Y-B^{*}Z\|^{2}\leq\min_{A,B}\mathbb{E}\|X-AY-BZ\|^{2}.$
10.4. In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors. Prove this fact as formally stated in Lemma 10.6. Conclud that the Kalman filter can be learned using Algorithm 6.
10.5. Prove Lemma 10.8, which is stated as follows. If a random variable pair $(\mathbf{x},\mathbf{y})$ are distributed as
then the conditional distribution of $\mathbf{x}$ given $\mathbf{y}$ can be written as $\mathbf{x}|\mathbf{y}\sim\mathcal{N}(\mu_{x}+\Sigma_{12}\Sigma_{22}^{-1}(\mathbf{y}-\mu_{y}),\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}).$
@@
Chapter 11 Spectral Filtering
In the previous chapter, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
  1. 1.
    State estimation via Kalman filtering, which recursively estimates both the latent system state and the projection matrices that map the system state to observations. Kalman filtering provides an optimal solution under Gaussian noise, and updates the state estimates over time. However, it requires an accurate model of system dynamics, including transition and observation matrices, which may be difficult to estimate, especially in adversarial or high-noise environments.
  2. 2.
    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification. Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs. While this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the system’s stability constant, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.
In this chapter, we introduce Spectral Filtering, an alternative approach that bypasses these limitations. Rather than relying on state estimation or direct system identification, spectral filtering constructs predictors in a transformed basis, which does not depend on the system’s stability properties. This method leverages spectral decompositions to efficiently extract the relevant dynamical structure, enabling more robust and scalable learning in both stable and unstable systems.
To illustrate the power of spectral filtering, we first present a one-dimensional case, which provides an intuitive understanding of how spectral methods can be applied to system identification and forecasting. We then extend the discussion to higher-dimensional settings, demonstrating how spectral predictors enable a structured and computationally efficient representation of system dynamics.
While this chapter primarily focuses on symmetric system matrices for clarity, the principles of spectral filtering extend to asymmetric and more general linear dynamical system settings. These extensions require additional mathematical tools, which we reference in the bibliographic section.